OpenStudy (anonymous):

find the intersection(s) y=|x-4|+3 and -x + 2y = 5

OpenStudy (lynfran):

y=|x-4|+3 and -x+2y=5 rewrite the 2nd equation we get y=x/2+5/2 now equating y=|x-4|+3 and y=x/2 +5/2 |x-4|+3=x/2+5/2 removing the absolute valve sign we get +(x-4)+3=x/2+5/2 ... and -(x-4)+3=x/2+5/2 now we solve each equations...